Consider a step-index fiber for which n1 = 1.475, n2 = 1.460 and a = 25µm. What is the maximum value of θ for which the ray ill be guided by the fiber. 


1
Expert's answer
2021-05-06T18:47:55-0400

"\sin \theta=\sqrt{n_1^2-n_2^2},"

"\sin\theta=\sqrt{1.475^2-1.46^2}=0.209,"

"\theta=12.1°."